Radium Ra236 has a half-life of 1590
years.an. How much of the original amount
of Ra236 would remain after 6360 year
?
 

Grade:12

2 Answers

Mahima Kanawat
5 years ago
Dear student....Radioactive reactions follow first order kinetics, so, N/N°=1/2^(t/T) ; 1/2^6360/1590 =1/2^4=1/16, therefore remaining amount would be 1/16 times of original amount... WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)
ADARSH SINGH
5 years ago
friend,
here goes your answer....
since, N=NO/2n where n=t/t1/2
here time (t) given is 6360 yr
and t1/2=1590 yr therefore, n=6360/1590 =4
on substituting this value we get,
N/No=1/2n= 1/24 =1/16
1/16 IS YOUR FINAL ANSWER...
 HOPE YOU GOT YOU ANSWER...
THANKU......
REPLY IF ANY QUERIES....